OpenStudy (anonymous):

The slope-intercept equation of the line through the points (3,7) and (6,4) is y=mx+b

OpenStudy (anonymous):

Slope= change in y/ change in x Slope = -3/3 = -1 y = -x+b Keeping that slope when x = 0, y = 10 y = -x + 10

OpenStudy (anonymous):

Thank you so much!